# Can industrial overcapacity enable seasonal flexibility in electricity use? A case study of aluminum smelting in China

### 26 Mar 2026 v2

@@ 4.7 Limitations and opportunities
Our study has several limitations that offer avenues for future research. Our model employs a top-down approach, projecting nationwide aluminum demand, which may differ from granular end-use forecasts [30]. However, our projected trends align with those of bottom-up studies [62]. We also simplify the distinction between primary/recycled aluminum, as well as market/trade dynamics. Regarding smelter representation, our model aggregates plants at the provincial level (see Supplementary Note 5.2 for details) and, due to computational constraints, assumes fixed capacity ratios between provinces. This spatial aggregation likely results in a conservative estimate of flexibility value, as it ignores facility-level modularity—in which individual potlines or plants could be shut down more efficiently than provincial-level ramping—and overlooks intra-provincial grid constraints and technological heterogeneity among plants. The fixed capacity ratio assumption further prevents us from deriving an optimized, facility-level retirement strategy. In reality, a more granular representation of retirement decisions would likely favor retaining plants that are more modern and technically efficient, which would be better equipped to provide grid flexibility. Future work could leverage efficient frameworks to optimize the spatial distribution of smelter retirement or migration, thereby capturing regional flexibility value, or, with enough data, offer plant-level operational strategies. We also do not differentiate between smelting technologies, retrofitting options, or stoichiometry of aluminum smelting [46, 53]. Primary aluminum is uniquely suited for seasonal stockpiling due to its homogeneity and negligible storage degradation, whereas other commodities (e.g., steel, cement) face higher physical risks and holding costs. When extending this framework to other heavy industries, future research must incorporate sector-specific storage degradation and holding costs. Finally, our co-optimization focuses on the aluminum smelting industry — the most electrified energy-intensive industry. Subsequent research should explore the co-optimization of multiple energy-intensive sectors to develop a more general understanding of how other energy-intensive industries can collectively manage structural overcapacity and navigate the low-carbon transition. More broadly, the question of how to leverage industrial-energy sector coupling to mitigate the socio-economic impacts of energy decarbonization and industrial restructuring represents a promising future research direction.
@@ Acknowledgment
The authors thank Tengmu Li for his valuable perspectives on the seasonal socio-economic rhythms of the contemporary workforce, Zongzhi Du for providing consultation on the operation of aluminum corporations in China, and Eric Larson for reviewing the paper.

### 17 Apr 2026 v3

@@ Methods
Our study has several limitations that offer avenues for future research. Regarding smelter representation, our model aggregates plants at the provincial level (see Supplementary Note 5.2 for details) and, due to computational constraints, assumes fixed capacity ratios between provinces. This spatial aggregation overlooks technological heterogeneity among plants. The same capacity retaining ratio across the nation prevents us from deriving an optimized, facility-level retirement strategy. In reality, a more granular representation of retirement decisions would likely favor retaining plants that are more modern and technically efficient, which would be better equipped to provide grid flexibility. Future work could leverage efficient frameworks to optimize the spatial distribution of smelter retirement or migration, thereby capturing regional flexibility value, or, with enough data, offer plant-level operational strategies. We also do not differentiate between smelting technologies, retrofitting options, or stoichiometry of aluminum smelting [47, 53]. Aluminum smelting is the most electrified energy-intensive industry and is uniquely suited for seasonal stockpiling due to its homogeneity and negligible storage degradation, whereas other commodities (e.g., steel, cement) face higher physical risks and holding costs. When extending this framework to other heavy industries, future research must incorporate sector-specific storage degradation and holding costs. Subsequent research could explore the co-optimization of multiple energy-intensive sectors to develop a more general understanding of how other energy-intensive industries can collectively manage overcapacity and navigate the low-carbon transition. More broadly, the question of how to leverage industrial-energy sector coupling to mitigate the socio-economic impacts of energy decarbonization and industrial restructuring represents a promising future research direction.
Data availability
All PyPSA input and results datasets relevant to this study are available via Zenodo at https://doi.org/10.5281/zenodo.19600477[36]. Additional data are available from the corresponding author on reasonable request.
Code availability
The PyPSA Power System Analysis model is open source and available via GitHub at https://github.com/PyPSA/PyPSA. Source code for the modified version of PyPSA-China used in this work is available via Github at https://github.com/Rick10119/PyPSA-China-THU.
@@ Acknowledgment
This work was supported by National Natural Science Foundation of China (No. 52130702, No. 72342007, No. 72422015) and Carbon Neutrality and Energy System Transformation project initiated by Tsinghua University. The simulations presented in this article were performed on the Della cluster at Princeton University, which is managed by the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s Research Computing department. The authors thank Tengmu Li for perspectives on the seasonal rhythms of the contemporary workforce, Zongzhi Du for providing consultation on aluminum corporations in China, and Eric Larson for providing feedback on the manuscript.
Author contributions
R.L., H.G., C.K. and J.D.J. conceptualized the study. R.L. and J.D.J. developed the experimental design. R.L., Y.S. and H.G. designed and performed the smelter simulations. R.L. developed the optimization model and other input datasets. R.L. performed the formal analysis, visualization and investigation, and produced the figures. R.L. drafted, revised and finalized the manuscript. J.D.J., A.L., H.L., H.G., J.W., E.D. and C.K. advised on the analysis and reviewed and revised the manuscript.
Competing interests
J.D.J. is a co-founder and Chief Technology Officer of Firma Power, LLC, which develops optimization and decision support software for energy systems. J.D.J. is part owner of DeSolve, LLC, which provides techno-economic analysis and decision support for clean energy technology ventures and investors. A list of clients can be found at https://www.linkedin.com/in/jessedjenkins. He serves on the advisory boards of Eavor Technologies Inc., a closed-loop geothermal technology company; Rondo Energy, a provider of high-temperature thermal energy storage and industrial decarbonization solutions; Dig Energy, a developer of low-cost drilling solutions for ground-source heat pumps; Karman Industries; and Emerald AI and has an equity interest in each company. He also serves as a technical advisor to MUUS Climate Partners and Energy Impact Partners, both investors in early-stage climate technology companies.